# Spatial Context Improves the Integration of Text with Remote Sensing for Mapping Environmental Variables

### v1

@@ Abstract
1. Recent developments in natural language processing highlight text as an emerging data source for ecology. Textual resources carry unique information that can be used in complementarity with geospatial data sources, such as remote sensing imagery, thus providing insights at the local scale into environmental conditions and properties hidden from more traditional data sources.
2. Leveraging textual information in a spatial context presents several challenges. First, the contribution of textual data remains poorly defined in an ecological context, and it is unclear for which tasks it should be incorporated. Unlike ubiquitous satellite imagery or environmental covariates, the availability of textual data is sparse and irregular; its integration with geospatial data is not straightforward. Moreover, text descriptions, typically those about environmental conditions, can be relevant for areas of variable size, sometimes over vast regions, introducing a high variability in the size of the spatial context to be integrated.
3. In response to these challenges, this work proposes an attention-based approach that combines aerial imagery and geolocated text within a spatial neighbourhood, i.e. integrating contributions from several nearby observations. Our approach combines vision and text representations with a geolocation encoding, allowing it to explicitly take into account the spatial structure of the original observations. The attention-based module dynamically selects spatial neighbours that are useful for predictive tasks, enabling the processing of a flexible range of inputs, with a spatial context of dynamic size.
4. The proposed approach is applied to the EcoWikiRS dataset, which combines high-resolution aerial imagery with sentences extracted from Wikipedia describing local environmental conditions across Switzerland. Our model is evaluated on the task of predicting 103 environmental variables from the SWECO25 data cube. Our approach consistently outperforms single-location or unimodal, i.e. image-only or text-only, baselines. When analysing variables by thematic groups, results show a significant improvement in performance for climatic, edaphic, population and land use/land cover variables, underscoring the benefit of including the spatial context when combining text and image data. Overall, our findings demonstrate an advantage of injecting knowledge through text when spatially modelling environmental variables with remote sensing imagery.
@@
Remote Sensing, Spatial Context, Spatial Interpolation, Switzerland, Text, Vision-Language Models, Wikipedia.
Figure 1: We study how to effectively combine remote sensing images and geolocated text for predicting a variety of environmental variables with a single deep learning model. For a location of interest ($x_{0}$), we use the geolocation of species observations in crowdsourcing platforms as a way to introduce geolocated text: We combine aerial images ($I_{0}$ to $I_{3}$) with text describing the species habitat extracted from its Wikipedia article ($T_{0}$ to $T_{3}$) and encode them with pretrained vision and text encoders, respectively. We incorporate several spatial neighbours that are in the vicinity of the location of interest and keep the spatial organisation of the observation by encoding geolocation ($x_{0}$ to $x_{3}$). We fuse all the textual and visual inputs through an attention-based fusion module that dynamically selects which neighbours are useful.
@@ 1 Introduction
Recent advances in machine learning applied to remote sensing images have enabled large-scale, cost-effective monitoring of the natural environment, supporting a variety of applications such as terrestrial plant diversity mapping (Wang and Gamon, 2019), canopy height estimation (Lang et al., 2023), species distribution modelling (Deneu et al., 2022), offshore industrial activities mapping (Paolo et al., 2024), population estimation (Metzger et al., 2022), bioclimatic variables regression (Waltari et al., 2014) or water quality monitoring (Sun et al., 2022).
The whole complexity of natural environments is rarely captured by a single data source. The use of heterogeneous data sources is often essential for a comprehensive understanding of ecological dynamics, with approaches that combine remote sensing with additional sources such as bioclimatic variables, in situ observations or laboratory measurements (Cavender-Bares et al., 2022; Miao et al., 2024; Liu et al., 2024; Abdelwahed et al., 2025; Zbinden et al., 2025). These heterogeneous data sources provide complementary perspectives on natural processes, each contributing unique information revealing details possibly not available in other sources. Integration of heterogeneous data sources can be challenging, particularly for those that are not available regularly, i.e. following a sparse sampling process, but recent approaches show first successes for this type of multimodal logic.
A new promising ecological data modality is text. The integration of textual information in environmental modelling pipelines holds promises and could bring significant contributions to applications at the intersection of natural language processing and ecology (Xue and Zhu, 2025; Miao et al., 2024). Of particular interest is the possibility of injecting knowledge through text into machine learning models. Automatic methods to process text enable information retrieval and structuring that would otherwise require several time-consuming, manual pre-processing steps (Castro et al., 2024). Various textual resources have been exploited to obtain specialised knowledge for a variety of applications, such as collecting plant traits systematically from the web (Marcos et al., 2025; Domazetoski et al., 2023; Paz-Argaman et al., 2020), monitoring illegal wildlife trade from online trade platforms (Stringham et al., 2021), or quantifying temporal ecological patterns from social media posts (Hart et al., 2018; Miłodrowski et al., 2025). The emergence of Large Language Models (LLMs) has also facilitated the filtering and rephrasing of the information coming from text, enabling an easier integration into machine learning models (Hamilton et al., 2024; Wen et al., 2025; Castro et al., 2024). Given the risks of LLMs to produce hallucinations and to introduce spurious biases in modelling pipelines, the relevance and correctness of the collected text content have also been assessed in recent works (Castro et al., 2024; Miao et al., 2024; Dorm et al., 2025).
The integration of text with other modalities has been significantly improved with Vision-Language Models (VLMs). CLIP (Radford et al., 2021), one well known VLMs, leverages the rich language semantics of image–text pairs extracted from the web for predicting image labels in zero-shot settings, i.e. predicting an image label without being explicitly trained for it. CLIP-like models have been widely leveraged: On the one hand, language semantics facilitate the transferability of annotations for real-world applications by harmonising different nomenclatures. One the other hand, the use of free-form text can be used as a supervision to introduce human expertise in a flexible manner in visual recognition tasks. For instance, WildCLIP (Gabeff et al., 2024) adapts the CLIP model to retrieve scenic, behavioural, and species attributes from camera trap images with user-defined queries expressed with language. BioCLIP (Gu et al., 2025; Stevens et al., 2024) leverages the natural hierarchy between species taxonomic names to improve fine-grained species recognition. INQUIRE (Vendrow et al., 2024) introduces a challenging text-to-image retrieval benchmark with textual descriptions that reflect real-world scientific queries, encouraging the development of retrieval systems that can support ecological and biodiversity research.
To study ecological geospatial processes, texts need to be associated with a geographic location. Geotagged social media posts (Häberle et al., 2022; Sathianarayanan et al., 2024; Yang et al., 2022), volunteered geographic information (VGI) such as OpenStreetMap (Wang et al., 2024; de Arruda et al., 2024; Fonte et al., 2018), or web content associated with geographic coordinates (Uzkent et al., 2019) are frequently used as a mean to associate a text with geographic coordinates. Alternatively, some texts can be associated with a geolocation by exploiting information explicitly present in the text (e.g. addresses, toponyms), or a relative description of the place, and also by identifying specific language features, such as dialects and idioms, to infer the location of the speaker (Derungs and Purves, 2014; Stock et al., 2022).
In ecology, a creative way to attach text to geographic coordinates is to exploit geo-tagged wildlife observations from crowd-sourcing platforms, where geolocations are paired with textual descriptions of the species (Daroya et al., 2025; Hamilton et al., 2024; Zermatten et al., 2025; Lange et al., 2025). Species descriptions, such as those coming from Wikipedia, provide valuable details on the ecological preferences of species, with clues on land cover, soil properties, relationship to humans, climatic conditions, distribution range and co-occurrence with other species. Thus, species observations are a rich source of information for ecology, since they reflect the local natural conditions around each observation, and are available globally. Combined with other resources, they have shown to be beneficial for several geospatial tasks. For instance, integrating text from Wikipedia in WildSAT (Daroya et al., 2025) has improved the recognition performance on satellite imagery at a global scale. The authors of EcoWikiRS (Zermatten et al., 2025) show that sentences describing local environmental conditions contribute to a better identification of fine-grained ecosystems and their characteristics from aerial imagery, while in (Hamilton et al., 2024), habitat or range descriptions enhance the accuracy of species distribution maps for species with few observations.
The use of text within geospatial ecological modelling is promising, yet faces significant obstacles. First, text is not a point measurement and exhibits geolocation ambiguity. Some texts, such as those describing natural features, are often valid over broad regions, and it is not always possible to link them to a specific area, such as municipal boundaries or standard land cover polygons. This geolocation ambiguity reflects not only the precision with which an ecological text is geolocated, but also the range for which a text is considered a valid descriptor. Instead of assigning text to a specific point or area, results can be shared as a probability distribution rather than a pair of coordinates (Dufour et al., 2025). Second is geographic biases. A large share of textual resources relies on crowd-sourced data, which are socio-economically and geographically biased towards the global north and populated places, while species observations are biased towards charismatic taxa (Wolf et al., 2022). Geo-referenced text often exhibits an uneven spatial distribution or an imprecise geolocation, resulting in noisy and inconsistently spread data. Such sparsity of textual data makes integration with continuous data (e.g., imagery, environmental rasters, etc.) challenging and demands specific integration methodologies.
The irregularity and sparsity of text data make it difficult to exploit methods relying on a fixed neighbourhood and tend to lead to inconsistent estimates. Transformations to continuous representations (interpolation) or density are often used instead. Methods relying on geostatistics have been developed to explicitly use spatial dependencies in interpolation: Kriging (Cressie, 1990) exploits spatial dependencies to characterise the range of autocorrelation and predict values at unobserved locations, and is widely used in disciplines such as meteorology, e.g. for snow depth prediction or temperature estimation. Geographically Weighted Regression (GWR) (Fotheringham et al., 2009) incorporates spatial dependencies by fitting local regression parameters, in which nearby samples exert greater influence on parameter estimates, and has been used to estimate the net primary production of forest ecosystems (Wang et al., 2005). Graph Neural Networks (GNNs) learn relevant geographic context by optimally weighting the spatial relationships across training samples. Recent works combine geostatistics with deep learning methods: Kriging Convolutional Networks (Appleby et al., 2020) merge the advantages of GNNs and Kriging by directly combining neighbours to make spatial predictions, for example, in bird count modelling or for estimating precipitation distribution. Similarly, Deep Kriging Neural Networks (Chen et al., 2024) include a deep-learning-based spatial encoder and a Kriging-based decoder and obtain a high performance for modelling meteorological data in the continental United States.
@@
Location encoding methods have also been explored to integrate geospatial context (Klemmer et al., 2025b). They encode geographic coordinates into a high-dimensional vector, or location embedding, that is directly usable by machine learning models. Numerous studies have shown the benefit of location encoding to improve model performance in many geospatial applications, ranging from species distribution modelling, fine-grained species recognition, or climate variables estimation (Mac Aodha et al., 2019; Cole et al., 2023; Chu et al., 2019; Rußwurm et al., 2024; Lange et al., 2025). Embeddings that capture the local characteristics of a place have also been built upon visual representations from satellite (Klemmer et al., 2025a) or ground-level imagery (Mai et al., 2023).
In this work, we study how to synergise text information from crowdsourced resources with very high-resolution remote sensing imagery to regress environmental and climatic variables. Xue and Zhu (2025) recently highlighted the potential of such synergy, which enables the prediction of continuous, real-valued environmental or geophysical variables (e.g. forest height, climatic variables, hydrology, population estimate). However, the absence of dedicated benchmarks and the challenge of combining discrete (text) and continuous (RS imagery or other grid-like data) variables were raised as core obstacles for the development of this field. Tackling these issues, we focus on three aspects: the contribution of text, strategies to integrate it, and a dataset for evaluation.
  • 1.
    First, we investigate whether text provides complementary information to remote sensing data for mapping environmental variables. To evaluate this both quantitatively and qualitatively, we compare the contributions of text and RS imagery in a multiple regression task predicting 103 environmental variables simultaneously. The variables are grouped into $7$ groups: vegetation, climate, geology, hydrology, soil (edaphic), population, and land use/land cover.
  • 2.
    Second, we address the challenge of combining sparsely geolocalised text with continuous RS imagery by proposing an attention-based architecture that explicitly takes into account the geospatial context of the observations. We propose to learn from a spatial context that goes beyond the strict location of a point of interest and leverages nearby observations, therefore improving robustness against noisy or sparse data. Practically, this is implemented as a form of spatial smoothing, by combining data at the point of interest with observations from the nearest neighbours in terms of geographical distance. We also preserve the spatial arrangement between samples with location embeddings. The attention-based architecture dynamically selects the most relevant neighbours by including a notion of geographic distance, while dynamically learning the range of the spatial context that is useful.
  • 3.
    We test our approach on the EcoWikiRS dataset (Zermatten et al., 2025) composed of aerial imagery over the Swiss territory and geo-referenced text extracted from Wikipedia that describes the local environmental context. As target, we use $103$ variables from SWECO25 (Külling et al., 2024), a raster database for ecological research, extracted at each location present in the EcoWikiRS dataset. We release this evaluation dataset, addressing the current lack of standardised comparison for VLMs in spatial regression tasks (Xue and Zhu, 2025).
We believe that the proposed approach sheds light on the role of text for geospatial regressions and has the potential to address challenges related to the incorporation of sparse texts with continuous RS imagery. Moreover, the approach is broadly transferable to other domains and tasks in ecology that integrate spatially geolocated textual information.
@@ 2 Materials and Methods
Figure 2: The proposed model architecture is illustrated with a spatial context of 3 nearest neighbours (k=3), with images, text and their respective geolocations as inputs. The pretrained encoders for image and text are frozen, while the location encoder and the attention-based fusion module are trained. Random token masking (dashed arrows) is applied on the generated visual ($v_{i}$) and textual ($t_{i,j}$) embeddings after the fusion with the location encoding ($LOC_{i}$). The final predictions are derived from the classification token (CLS) at the output of the attention-based fusion module. Colours denote distinct geographic locations.
The proposed approach integrates geolocated text and aerial images from nearby locations to predict 103 environmental variables using a multiple regression approach. Our multimodal approach is based on a vision-language transformer model that is able to ingest both remote sensing images and textual descriptions. The overall model architecture is shown in Fig. 2. We adopted a modular architecture that consists of an image encoder, a text encoder, a location encoder, and an attention-based fusion module. Similar to a graph-based approach or to GWR, the model learns to estimate a different weight for each nearby observation at each location of interest through its attention mechanism.

### v2

# A Geolocation-Aware Multimodal Approach for Ecological Prediction

@@ Abstract
1. Ecological monitoring increasingly relies on heterogeneous data sources, including remote sensing imagery, biodiversity observations, climatic variables, or textual knowledge. While integrating multiple modalities has the potential to improve environmental prediction, current approaches struggle to combine data sources with heterogeneous formats or contents. A central difficulty arises when combining continuous gridded data (e.g., remote sensing) with sparse and irregular point observations such as species records. Existing geostatistical and deep-learning-based approaches typically operate on a single modality or focus on spatially aligned inputs, and thus cannot seamlessly overcome this difficulty.
2. In response to these challenges, we propose a Geolocation-Aware MultiModal Approach (GAMMA), a transformer-based fusion approach designed to integrate heterogeneous ecological data using explicit spatial context. Instead of interpolating observations into a common grid, GAMMA first represents all inputs as location-aware embeddings that preserve spatial relationships between samples. Then, through attention mechanisms, GAMMA dynamically selects relevant neighbours across modalities and spatial scales, enabling the model to jointly exploit continuous remote sensing imagery and sparse geolocated observations.
3. We evaluate GAMMA on the task of predicting 103 environmental variables from the SWECO25 data cube across Switzerland. Inputs combine aerial imagery with biodiversity observations from GBIF and textual habitat descriptions from Wikipedia, provided by the EcoWikiRS dataset.
4. Experiments show that multimodal fusion consistently improves prediction performance over single-modality baselines and that explicit spatial context further enhances model accuracy. The flexible architecture of GAMMA also allows to analyse the contribution of each modality through controlled ablation experiments. These results demonstrate the potential of location-aware multimodal learning for integrating heterogeneous ecological data and for supporting large-scale environmental mapping tasks and biodiversity monitoring. The source code is included in the submission files.
@@
Multimodal Learning, Remote Sensing, Spatial Context, Text, Vision-Language Models.
Figure 1: Overview of the proposed Geolocation-Aware MultiModal Approach (GAMMA). Ecological data assimilation pipelines are often based on single-point observations (a) or on nearby records within a local fixed neighbourhood (b), as commonly exploited in CNNs or GNNs. Leveraging both nearby and distant observations across modalities (c) enables capturing broader spatial dependencies. The proposed GAMMA (d) addresses this limitation by representing observations as location-aware embeddings and using a transformer encoder to dynamically select the most relevant neighbours across modalities and spatial scales. GAMMA effectively combines remote sensing imagery and sparse geolocated text for predicting diverse environmental variables within a single model.
@@ 1 Introduction
Modern ecological and environmental research increasingly relies on heterogeneous spatial datasets: climatic variables, camera trap photographs, satellite imagery, citizen science records, tabular data, audio, in situ measurements and even text information (Cavender-Bares et al., 2022; Liu et al., 2024; Pollock et al., 2025). Each modality provides a distinct, but complementary perspective on natural processes, revealing details possibly not available in other sources. Integrating these diverse modalities into a unified framework is widely recognised to increase predictive performance. However, the selection of an appropriate data fusion approach remains a central challenge in data assimilation (Isaac et al., 2020; Zipkin et al., 2021; Tuia et al., 2022; Miao et al., 2026).
A major difficulty in ecological data assimilation arises from their interoperability: the integration of heterogeneous data sources, e.g. continuous and discrete observations, remains a challenge (Martinez et al., 2025; Jung, 2023; Koldasbayeva et al., 2024; Wicquart et al., 2022). Continuous sources, such as remote sensing imagery or climatic variables, carry information at every location in a domain and are typically provided as regularly gridded raster data, facilitating spatial alignment across variables. In contrast, many ecological observations are recorded as discrete measurements tied to specific geographic coordinates. Examples include field measurements and species surveys, which are often irregularly distributed in space and describe properties valid over a limited (and often variable) spatial footprint. Combining these heterogeneous spatial representations requires methods capable of handling irregular sampling, exploiting spatial autocorrelation, and propagating information across space.
Addressing this challenge has been a long-standing objective in spatial statistics. Geostatistical methods have historically addressed this problem by interpolating sparse observations into continuous spatial fields. Kriging (Cressie, 1990; Matheron, 1967) or Geographically Weighted Regression (GWR) (Fotheringham et al., 2009) use spatial autocorrelation to generate predictions at unobserved locations and have long been central tools for environmental (e.g. biomass (Wang et al., 2005)) and geological mapping (Song et al., 2016), as well as in meteorology (e.g. for snow depth prediction or temperature estimation (Cressie, 1990)).
Recent advances in deep learning have revisited these ideas by learning spatial relationships directly from data. Graph Neural Networks (GNNs) (Wu et al., 2020) explicitly model relationships between observations and can learn relevant geographic context by optimally weighting the spatial relationships across training samples. These approaches have been applied to several geospatial tasks, including land use classification (Xu et al., 2022) or urban environment analysis (Zhu et al., 2020). Building on this idea, several studies combine geostatistics with deep learning methods; Kriging Convolutional Networks (Appleby et al., 2020), for instance, merge the advantages of GNNs and Kriging by directly combining neighbours for spatial prediction tasks, such as bird count modelling or precipitation estimation. Similarly, Deep Kriging Neural Networks (Chen et al., 2024) include a deep learning-based spatial encoder and a Kriging-based decoder for modelling meteorological data. Tian and Shan (2025) exploits the Vision Transformer (Dosovitskiy et al., 2021) (ViT) architecture to interpolate a digital elevation model from sparse altimetry data. Overall, these methods substantially improve predictive performance over traditional approaches, but remain largely unimodal.
In parallel to approaches designed for sparse observations, spatial context in regularly gridded data has been extensively exploited through convolutional neural networks (CNNs). CNNs naturally encode local spatial neighbourhoods by applying convolutional filters over raster grids, allowing models to learn spatial patterns directly from neighbouring pixels. This property has made CNNs a dominant approach in remote sensing tasks such as land cover classification (Tuia et al., 2024), population mapping (Metzger et al., 2022) or vegetation monitoring (Lang et al., 2023). ViTs have extended this idea by modelling spatial context through self-attention over image patches, enabling the integration of larger spatial contexts while preserving the grid structure of remote sensing imagery (Lu et al., 2025; Xiao et al., 2025). Overall, these approaches typically address either dense raster data or sparse point observations, but not both simultaneously.
@@
Instead of interpolating geospatial information from observations, an alternative line of work incorporates spatial information directly into the model through location encoding. Location encoding methods transform geographic coordinates into high-dimensional embeddings that capture spatial relationships between locations (Mac Aodha et al., 2019; Mai et al., 2022; Cole et al., 2023; Rußwurm et al., 2024). Beyond geographic proximity, these embeddings can also incorporate contextual information derived from auxiliary data such as satellite imagery (Klemmer et al., 2025), ground-level imagery (Mai et al., 2023) or climatic variables (Daroya et al., 2025). Location embeddings are beneficial in many geospatial applications, from species distribution modelling, fine-grained species recognition, to climate variables estimation (Cole et al., 2023; Chu et al., 2019; Lange et al., 2025).
Transformers (Vaswani et al., 2017), which excel at integrating heterogeneous sequences of inputs, offer a natural solution to integrate multiple data modalities. Its architecture has proven highly effective in vision-language models such as VisualBERT (Li et al., 2020) or LLaVA (Liu et al., 2023), which process image and text representations jointly, enabling rich cross-modal reasoning. This same principle can be applied to geospatial data, where heterogeneous inputs can be treated as complementary representations describing a shared context.
Two recent approaches leveraging transformers for geospatial prediction are particularly related to our work. Space2Vec (Mai et al., 2020) uses a transformer architecture to incorporate location embeddings of multiple geographic objects to predict Points of Interest (POIs) type or image category based on their geolocation. However, similarly to Kriging, Space2Vec uses neighbouring target values as inputs (e.g. POI types of neighbouring locations), restricting its applicability to settings where such labels are not available. Lange et al. (2025) presents an attention-based approach to predict species range maps based on multiple geolocated observations in a few-shot setting. This approach is primarily coordinate-centric, without exploring relative positional encodings, and adding, optionally, one non-geocoded sentence or image as metadata. Addressing these limitations calls for a framework that requires no neighbour supervision, is inherently multimodal, and explicitly encodes spatial relationships across heterogeneous input sources.
We propose a Geolocation-Aware MultiModal Approach (GAMMA) that integrates multimodal ecological information sources using explicit location encoding. As illustrated in Fig. 1, GAMMA learns spatial context beyond a single location (1(a)) or the fixed local neighbourhood typically captured by CNNs (1(b)). Instead, it leverages sparse observations and gridded data within an arbitrarily large neighbourhood (1(c)). This is implemented as location embeddings that preserve the spatial arrangement between samples. Rather than interpolating discrete sources into a continuous spatial field, GAMMA represents all inputs as location-aware representations and uses a transformer encoder to learn cross-modal spatial relationships directly from data. Similar to a graph-based approach or to GWR, GAMMA dynamically selects the most relevant neighbours by including a notion of geographic distance, and fuses them according to feature similarity. Furthermore, GAMMA naturally accommodates flexible input combinations, enabling systematic ablation of individual modalities to quantify the independent contribution of each data source.
We evaluate GAMMA on the prediction of 103 environmental variables from the SWECO25 (Külling et al., 2024) data cube. SWECO25 variables span hydrological, climatic, geological, vegetation, population, land cover, and soil properties, and together form a comprehensive characterisation of a location that no single sensor can capture alone. As input modalities, we integrate remote sensing imagery as gridded variables with biodiversity observations from GBIF and textual habitat descriptions from Wikipedia, sourced from the EcoWikiRS (Zermatten et al., 2025) dataset. Similar to recent works (Daroya et al., 2025; Lange et al., 2025; Hamilton et al., 2024), the species occurrence records serve as a means to assign a geolocation to Wikipedia-derived habitat sentences. Text provides insight into the local environmental conditions, which are potentially unavailable from the remote sensing perspective. SWECO variables are extracted at the locations corresponding to EcoWikiRS samples, forming a benchmark dataset that we release to address the current lack of benchmarks for vision-language models in geospatial regression tasks (Xue and Zhu, 2025).
The flexible input structure of GAMMA enables us to quantify the independent contribution of each modality, which is rarely systematically studied in vision and language models for ecological applications (Daroya et al., 2025; Hamilton et al., 2024; Lange et al., 2025; Miao et al., 2026; Xue and Zhu, 2025; Gu et al., 2025; Stevens et al., 2024; Wen et al., 2025). Here, we evaluate text-only, image-only, and combined settings, isolating what ecological knowledge each source provides. Our experiments demonstrate that multimodal fusion outperforms single-modality baselines, and explicit spatial context encoding yields additional synergistic gains over location-agnostic approaches. The proposed approach has the potential to address challenges related to the incorporation of sparse records with gridded data.
While focusing on vision and language inputs, GAMMA is broadly transferable to other domains and types of geospatial inputs in ecology or beyond. To summarise, this work makes three contributions:
  • We propose GAMMA, a geolocation-aware multimodal approach for integrating sparse and continuous spatial data.
  • We release an open benchmark dataset of 103 environmental variables combining EcoWikiRS and SWECO25 for spatial regression.
  • We provide a systematic evaluation of the contribution of textual ecological knowledge relative to remote sensing imagery.
@@ 2 Materials and Methods
Through the predictions of 103 SWECO environmental variables, GAMMA solves a multiple regression task by integrating sparse geolocated text and aerial images. The overall model architecture is shown in Fig. 2. We adopted a modular architecture that consists of an image encoder, a text encoder, a location encoder, and a transformer-based fusion module.
Figure 2: The proposed model architecture is illustrated with a spatial context of 3 nearest neighbours (k=3), with images, text and their respective geolocations as inputs. The pretrained encoders for image and text are frozen, while the location encoder and the transformer-based fusion module are trained. Random token masking (dashed arrows) is applied on the generated visual ($v_{i}$) and textual ($t_{i,j}$) embeddings after the fusion with the location encoding ($LOC_{i}$). The final predictions are derived from the classification token (CLS) at the output of the transformer module. Colours for $LOC_{i}$ tokens denote distinct geographic locations.